Given y is two times x, z is three times x, and x + y + 3z = 12. What are the values of x, y, and z ?
Answer by stanbon(75887)

Given y is two times x, z is three times x, and x + y + 3z = 12. What are the values of x, y, and z ?
----
Equations:
y = 2x
z = 3x
x + y + 3z = 12
----
Substitutue and solve for "x":
x + 2x + 9x = 12
12x = 12
x = 1
---
y = 2x = 2
z = 3x = 3
============================
Cheers,
Stan H.
=============